OpenStudy (anonymous):

with graph y=x2+16 solve x2-16=0

OpenStudy (anonymous):

I am not sure what the equation y=x^2+16 has anytihng to do with solving for x^2-16... solving for x^2-16=0 however canbe done fairly simply using algebra x^2-16=0 First add 16 to both sides x^2=16 then take the squareroot of both sides \[x=\pm4\] and that's your solution. Let me know if I misread the question.

OpenStudy (anonymous):

thanks, i not sure what relevance the graph had ?

OpenStudy (anonymous):

I'm not exactly sure either.